Question 8 (4 points) anagement must decide between a full-price service using the companys new fleet of jet aircraft and a discount service using smaller-capacity commuter planes. It is clear that the best choice depends on the market reaction to the service Myrtle Air offers. Management developed estimates of the contribution to profit for each type of service based on three possible levels of demand for service to Myrtle Beach: strong, medium and weak. The following table shows the estimated quarterly profits (in thousands of dollars): Demand for Service Service Strong Medium Weak Full Price 950 Discount 750 450 Probability 0.4 0.3 600 -400 300 0.3 Suppose that management of Myrtle Air Express has developed the following utility function: U(950) 10o, U(750) 87, U(600) 72, U(450) - 58, U(300) - 48, and U(-400) 0 Use the expected utility approach to determine the optimal decision. What is the expected utility of the optimal choice? (Round to one decimal place) Your Answer: Answer Save
Answer #1

for full price
expected utility = 0.4 U(950) + 0.3 U(600) + 0.3 U(-400)
= 0.4 * 100 + 0.3 * 72 + 0.3 * 0
= 40 + 21.6
=61.6
for Discount
expected utility
= 0.4 *U(750) + 0.3*U(450) + 0.3 * U(300)
= 0.4 * 87+ 0.3 * 58 + 0.3 * 48
= 66.6
since 66.6 > 61.6
we choose Discount and expected utility = 66.6
Please rate